# Capturing strong correlation effects on a quantum annealer: calculation of avoided crossing in the H4 molecule using the quantum annealer eigensolver

### v1

@@ V Conclusion and future outlook
In summary, we carry out calculations to predict an avoided crossing in the H4 molecule on a D-Wave quantum computer. The motivation for this work is two-fold: first, to address the challenge of capturing strong correlation effects, which are widely regarded as a cornerstone of physical chemistry, and second, to bridge a significant gap in literature with regard to the application of quantum annealers to NISQ-era chemistry, in contrast to the extensive body of research dedicated to the gate-based VQE algorithm in the same domain. We find from our calculations that using the QAE algorithm with 1000 shots and 10 repetitions yields AC with reasonable precision, to within about 20 percentage of the reference full configuration interaction values. Furthermore, we find that the result can be improved by increasing the number of shots to 5000, the number of repetitions to 30, and by choosing different anneal times for the ground and excited state energies. The choice of $\lambda$ range using SAE was sufficient. The improved AC result with 5000 shots and 30 repetitions is to within about 1.1 percent of the FCI value. Among the factors considered, we find that unsurprisingly, the number of shots significantly influences the quality of our results. It is also worth noting at this point that there are several other factors that could be tuned to improve the results further, which we do not carry out in this pilot study. These include, but are not limited to:
@@
  • Checking the effect of the number of physical qubits that are expended for a computation via embedding. We find that for a given repetition, that is, across hundred $\lambda$ values that are scanned for our chosen $\lambda$ range, the number of physical qubits can vary between about 700 to about 900 to construct our QUBO with 80 logical qubits. It is important to note that embedding is a heuristic, and it is worth exploring this direction further to improve the quality of results.
  • A study of the dependence of results on $K$. This is fairly non-trivial in practice, since for a Hamiltonian matrix of size $N$, an increase in $K$ (to achieve better precision in the CI coefficient values) leads to corresponding all-to-all connected QUBO growing as $NK$. While embedding larger graphs would likely incur more error, an acute problem that one would face is running out of physical qubits. For example, were we to carry out the same computations on H4 with a split valence basis set such as the 6-31G (4 electron- 5 orbital active space), a QUBO size of 170 (assuming $K=10$) requires on an average about 4000 physical qubits (over those scanned $\lambda$ values for which an embedding was found). Finding an embedding with any further increase in $K$ is not possible, since the D-Wave annealers have only 5000 physical qubits. In such cases, one needs to opt for Sub-QUBOs (for example, see Ref. [45]), and additional errors incurred in approximations involved in choice of sub-QUBOs also add to the error budget. It is worth adding that the embedding issue we consider is specific to current state-of-the-art D-Wave hardware and not to the QAE algorithm itself; future advances in this direction, for example, machines with substantially better connectivity, could alleviate embedding-related issues.
@@
We also carry out a qualitative comparison of the QAE algorithm with the celebrated VQE approach, and find that while both methods have their own pros and cons, QAE is a compelling NISQ-era contender for quantum chemistry on quantum computers.
@@ Acknowledgements.
This work was supported by the European Commission EIC-Transition project RoCCQeT (GA 101112839). AAZ and VSP sincerely acknowledge Palak Chawla for her tremendous support and patience in helping us with the VQE computations, and Disha Shetty for patiently reading the initial drafts and giving her feedback, which in turn helped improve the quality of the manuscript. VSP and AAZ acknowledge Prof. Bhanu Pratap Das for discussions on QAE and support with D-Wave time, and Prof. Debashis Mukherjee for discussions on the general nature of AC. The authors acknowledge Mr. Jose Miralles for fruitful discussions on Sub-QUBOs. KS acknowledges support from Quantum Leap Flagship Program (Grant No. JPMXS0120319794) from the MEXT, Japan, Center of Innovations for Sustainable Quantum AI (JPMJPF2221) from JST, Japan, and Grants-in-Aid for Scientific Research C (21K03407) and for Transformative Research Area B (23H03819) from JSPS, Japan.

### v4

@@ V Conclusion and future outlook
In summary, we have explored the potential of the QAE algorithm as a NISQ tool for carrying out electronic structure calculations by (a) predicting an avoided crossing in the H4 molecule on a D-Wave quantum computer, and (b) carrying out a preliminary scaling analysis comparison between QAE and the celebrated gate-based VQE algorithm. We anticipate that our analysis would serve as a starting point to bridge a significant gap in literature with regard to the application of quantum annealers to NISQ-era chemistry, in contrast to the extensive body of research dedicated to the gate-based VQE algorithm in the same domain.
We find from our calculations that using the QAE algorithm with 1000 shots and 10 repetitions yields AC with reasonable precision, to within about 20 percent of the reference full configuration interaction values. Furthermore, we find that the result can be substantially improved by increasing the number of shots to 5000, the number of repetitions to 30, and by choosing different anneal times for the ground and excited state energies. The choice of $\lambda$ range using SAE was found to be sufficient. The improved AC result with 5000 shots and 30 repetitions is to within about 1.2 percent of the FCI value. Among the factors considered, we find that unsurprisingly, the number of shots significantly influences the quality of our results. It is also worth noting at this point that there are several other factors that could be tuned to improve the results further, which we do not carry out in this pilot study. These include, but are not limited to:
@@
  • Studying the effect of the number of physical qubits that are expended for a computation via embedding.
@@
Via our scaling analysis, where we estimate costs for each step of the QAE and the VQE algorithms starting from problem inputs (HF, common to QAE and VQE), AO to MO integral transformation (common), CISD Hamiltonian matrix elements evaluation (specific to QAE), encoding the functional in its QUBO form (QAE), and input reference state isometry (VQE)) and then moving to the main processing steps (parameter update (VQE), convergence (VQE), number of two-qubit gates in the UCCSD ansatz circuit (VQE), embedding (VQE and QAE)), we find that performing CI calculations using QAE is a compelling NISQ-era contender for quantum chemistry on current-day quantum computers. Although both QAE and VQE scale worse relative to established classical methods (even under the strong assumption that an oracle supplies the input state for VQE), we find that they scale comparably between each other (it is also important to add that on grounds of the problem being very non-trivial, we do not estimate the scaling of the anneal time and number of required shots itself in QAE). We add that our conclusion comes with two caveats: (a) our analysis is to be seen as preliminary with scope for further refinement, and (b) we do not consider variants of VQE or QAE but rather their vanilla versions. Our hardware results indicate that in order to use the VQE algorithm to predict energies accurately within a considered active space, one requires quantum devices with gate fidelities well beyond the current best capabilities. On the other hand, the QAE algorithm is well-placed to obtain energies with reasonable precision on current-day D-Wave hardware. However, the next natural step of studying the effect of D-Wave hardware errors as we scale up the size of the molecular systems would require more physical qubits, in view of the embedding overheads. An alternative that has not been explored in this work is the use of sub-QUBOs to extend the calculations to larger system sizes in current D-Wave computers.
@@ Acknowledgements.
This work was supported by the European Commission EIC-Transition project RoCCQeT (GA 101112839). VSP acknowledges support from CRG grant (CRG/2023/002558) for some of the computations. AAZ and VSP sincerely acknowledge Palak Chawla for her tremendous support and patience in helping us with the VQE computations, and Disha Shetty for patiently reading the initial drafts and giving her feedback, which in turn helped improve the quality of the manuscript. VSP and AAZ acknowledge Prof. Bhanu Pratap Das for discussions on QAE and support with D-Wave time, and Prof. Debashis Mukherjee for discussions on the general nature of AC. The authors acknowledge Mr. Jose Miralles for fruitful discussions on Sub-QUBOs. KS acknowledges support from Quantum Leap Flagship Program (Grant No. JPMXS0120319794) from the MEXT, Japan, Center of Innovations for Sustainable Quantum AI (JPMJPF2221) from JST, Japan, and Grants-in-Aid for Scientific Research C (21K03407) and for Transformative Research Area B (23H03819) from JSPS, Japan. J.N and J.R thankfully acknowledge RES resources provided by Barcelona Supercomputing Center in Marenostrum 5 to FI-2025-1-0043.
VI Data availability
All the data is included in the article. The details of our code will be available on reasonable request.
Appendix A Derivation of fidelity bounds
The energy of a quantum state |$\ket{\Psi}$ is given by
$E=\langle\Psi|H|\Psi\rangle.$ (10)
Substituting the Hamiltonian, $H$, with its spectral decomposition, we obtain
$E$ $=$ $\sum_{n=0}^{d-1}E_{n}p_{n}$ (11)
$=$ $E_{0}p_{0}+\sum_{n=1}^{d-1}E_{n}p_{n},$
where $E_{n}$ is the energy of the eigenstate |$\ket{E_{n}}$, and $p_{n}=|\langle E_{n}|\Psi\rangle|^{2}$ represents its population.
Since the eigenvalues, $E_{n}$, are ordered in increasing magnitude, any energy, $E_{n}$, with $n\in{1,\cdots,d-1}$, satisfies the bounds:
$E_{1}\leq E_{n}\leq E_{max},$ (12)
where $E_{max}=E_{d-1}$ denotes the largest eigenvalue of $H$.
Table 3: Table showing worst-case scaling for the cost of computing different CISD Hamiltonian matrix elements. In the table, ($p,q,r,\ldots$ ) refer to general indices, ($i,j,k,\ldots$ ) refer to indices of occupied spin orbitals, and ($a,b,c,\ldots$) refer to indices of unoccupied spin orbitals.
S. No. |$\ket{\Phi_{m}}$ |$\ket{\Phi_{n}}$#\## same indices Examples of |$\ket{\Phi_{n}}$ Pattern Scaling
1. |$\ket{0}$ |$\ket{0}$ 0 |$\ket{0}$ 1 $\mathcal{O}(1)\times\mathcal{O}(1)\times\mathcal{O}(N^{2})=\mathcal{O}(N^{2})$
2. |$\ket{0}$ |$\ket{S}$ 0 |$\ket{\Phi_{i}^{a}}$ 2 $\mathcal{O}(1)\times\mathcal{O}(N^{2})\times\mathcal{O}(N)=\mathcal{O}(N^{3})$
3. |$\ket{0}$ |$\ket{D}$ 0 |$\ket{\Phi_{ij}^{ab}}$ 3 $\mathcal{O}(1)\times\mathcal{O}(N^{4})\times\mathcal{O}(1)=\mathcal{O}(N^{4})$
4. |$\ket{\Phi_{i}^{a}}$ |$\ket{S}$ 2 |$\ket{\Phi_{i}^{a}}$ 1 $\mathcal{O}(N^{2})\times\mathcal{O}(1)\times\mathcal{O}(N^{2})=\mathcal{O}(N^{4})$
5. |$\ket{\Phi_{i}^{a}}$ |$\ket{S}$ 1 |$\ket{\Phi_{i}^{b}}$, |$\ket{\Phi_{j}^{a}}$ 2 $\mathcal{O}(N^{2})\times\mathcal{O}(N)\times\mathcal{O}(N)=\mathcal{O}(N^{5})$
6. |$\ket{\Phi_{i}^{a}}$ |$\ket{S}$ 0 |$\ket{\Phi_{j}^{b}}$ 3 $\mathcal{O}(N^{2})\times\mathcal{O}(N^{2})\times\mathcal{O}(1)=\mathcal{O}(N^{4})$
7. |$\ket{\Phi_{i}^{a}}$ |$\ket{D}$ 2 |$\ket{\Phi_{ij}^{ab}}$, |$\ket{\Phi_{ij}^{ba}}$ 2 $\mathcal{O}(N^{2})\times\mathcal{O}(N^{2})\times\mathcal{O}(N)=\mathcal{O}(N^{5})$
8. |$\ket{\Phi_{i}^{a}}$ |$\ket{D}$ 1 |$\ket{\Phi_{ij}^{bc}}$, |$\ket{\Phi_{jk}^{ab}}$ 3 $\mathcal{O}(N^{2})\times\mathcal{O}(N^{3})\times\mathcal{O}(1)=\mathcal{O}(N^{5})$
9. |$\ket{\Phi_{i}^{a}}$ |$\ket{D}$ 0 |$\ket{\Phi_{jk}^{bc}}$ 4 -
10. |$\ket{\Phi_{ij}^{ab}}$ |$\ket{D}$ 4 |$\ket{\Phi_{ij}^{ab}}$ 1 $\mathcal{O}(N^{4})\times\mathcal{O}(1)\times\mathcal{O}(N^{2})=\mathcal{O}(N^{6})$
11. |$\ket{\Phi_{ij}^{ab}}$ |$\ket{D}$ 3 |$\ket{\Phi_{ik}^{ab}}$, |$\ket{\Phi_{kj}^{ab}}$, |$\ket{\Phi_{ij}^{ac}}$, |$\ket{\Phi_{ij}^{cb}}$, $\cdots$ 2 $\mathcal{O}(N^{4})\times\mathcal{O}(N)\times\mathcal{O}(N)=\mathcal{O}(N^{6})$
12. |$\ket{\Phi_{ij}^{ab}}$ |$\ket{D}$ 2 |$\ket{\Phi_{ij}^{cd}}$, |$\ket{\Phi_{ik}^{ac}}$, |$\ket{\Phi_{kj}^{ac}}$, |$\ket{\Phi_{kj}^{cb}}$, $\cdots$ 3 $\mathcal{O}(N^{4})\times\mathcal{O}(N^{2})\times\mathcal{O}(1)=\mathcal{O}(N^{6})$
13. |$\ket{\Phi_{ij}^{ab}}$ |$\ket{D}$ 1 |$\ket{\Phi_{ik}^{ab}}$, |$\ket{\Phi_{ik}^{cd}}$, |$\ket{\Phi_{kl}^{cb}}$, |$\ket{\Phi_{kl}^{ac}}$, $\cdots$ 4 -
14. |$\ket{\Phi_{ij}^{ab}}$ |$\ket{D}$ 0 |$\ket{\Phi_{kl}^{cd}}$ 4 -
Thus, the second term in the right hand side of Eq. (A2) is bounded as
$E_{1}(1-p_{0})\leq\sum_{n=1}^{d-1}E_{n}p_{n}\leq E_{max}(1-p_{0}),$ (13)
where we have used the normalization condition $\sum_{n=1}^{d-1}p_{n}=1-p_{0}$. Substituting this bound into Eq. (A2), the energy of the state |$\ket{\Psi}$ is bounded by
$E_{0}p_{0}+E_{1}(1-p_{0})\leq E\leq E_{0}p_{0}+E_{max}(1-p_{0}).$ (14)
After some algebra and identifying $p_{0}=|\langle E_{0}|\Psi\rangle|^{2}=\mathcal{F}$ as the fidelity of |$\ket{\Psi}$ with the ground state, we derive the following bounds for the fidelity:
$1-\frac{E-E_{0}}{E_{1}-E_{0}}\leq\mathcal{F}\leq 1-\frac{E-E_{0}}{E_{max}-E_{0}}.$ (15)
In our case, we check if the bounds are satisfied for the case of our avoided crossing problem for H4 in square geometry where $E_{0}=E_{g}$, $E_{1}=E_{e}$ and $E=E_{g,QAE}$.
The lower bound represents the case where the population outside the ground state is entirely in the first excited state, the one with the lowest energy. On the other hand, the upper bound corresponds to the scenario where all the population outside the ground state is in the state with the highest energy. In any realistic situation, the true population distribution will lie somewhere between these two extremes.
Appendix B Cost of evaluating CISD Hamiltonian matrix elements
Here, we present the cost of generating the CISD Hamiltonian matrix elements, $\langle\Phi_{m}|H|\Phi_{n}\rangle$, which are fed as inputs to QAE. The cost is analysed by categorizing the matrix elements into 4 patterns and applying the Slater–Condon rules. We begin by classifying $\langle\Phi_{m}|H|\Phi_{n}\rangle$:
  • Pattern 1: |$\ket{\Psi_{m}}=\ket{\Psi_{n}}$ (diagonal element of the CI matrix), for which the cost of evaluating the matrix element goes as $\mathcal{O}(N^{2})$.
  • Pattern 2: |$\ket{\Phi_{m}}$ and |$\ket{\Phi_{n}}$ have two different occupation indices (for example, |$\ket{\Phi_{m}}=\ket{\Phi_{i}^{a}}$ and |$\ket{\Phi_{n}}=\ket{\Phi_{i}^{b}}$ The cost of evaluating $\langle\Phi_{m}|H|\Phi_{n}\rangle$ then scales as $\mathcal{O}(N)$.
  • Pattern 3: |$\ket{\Psi_{m}}$ and |$\ket{\Psi_{n}}$ have four different occupation indices (for instance, |$\ket{\Phi_{m}}=\ket{\Phi_{i}^{a}}$ and |$\ket{\Phi_{ij}^{bc}}$). The cost of evaluating $\langle\Phi_{m}|H|\Phi_{n}\rangle$ then scales as $\mathcal{O}(1)$.
  • Pattern 4: |$\ket{\Psi_{m}}$ and |$\ket{\Psi_{n}}$ have more than four differing occupation indices (for instance, |$\ket{\Phi_{m}}=\ket{\Phi_{i}^{a}}$ and |$\ket{\Phi_{jk}^{bc}}$). These matrix elements are zero due to the Slater–Condon rules.
From the table, we see that the cost of evaluating the matrix elements for determinant-based CISD is at most $N^{6}$. We comment on why the scaling of serial number 5 is not $\mathcal{O}(N^{2})\times\mathcal{O}(N^{2})\times\mathcal{O}(N)=\mathcal{O}(N^{5})$. This is because one of the indices of |$\ket{\Phi_{m}}=\ket{\Phi_{j}^{b}}$ is fixed to either $j=i$ or $a=b$. Thus, the number of possible |$\ket{\Phi_{n}}$ scales as ${}^{2}C_{1}\times\mathcal{O}(N)=\mathcal{O}(N)$. Similarly, for serial number 11, the scaling is not $\mathcal{O}(N^{4})\times\mathcal{O}(N^{4})\times\mathcal{O}(N)=\mathcal{O}(N^{9})$, because three out of four indices of |$\ket{\Phi_{n}}$ should be fixed, and the number of possible |$\ket{\Phi_{n}}$ scales as ${}^{4}C_{3}\times\mathcal{O}(N)=\mathcal{O}(N)$.
In the case of GUGA-based CISD, we want to know the cost scaling for evaluating $\langle\Phi_{m}|H|\Phi_{n}\rangle$, where |$\ket{\Phi_{m}}$ and |$\ket{\Phi_{n}}$ are symmetry-adapted CSFs. The scaling behavior of the matrix elements evaluation is very similar to that of determinant-based one. However, in the GUGA-CI approach, we incur extra cost for evaluating an additional coefficient depending on the pathways of Shavitt graph, with the cost for this step possibly scaling as $\mathcal{O}(N)$.
Appendix C Details of input state preparation: a representative example
Here, we present the details of the conditions under which our representative example yielded the $2^{N}$ scaling for the number of two-qubit gates for the isometry associated with input state preparation for VQE. We consider a simplified model of a two-electron system where the number of determinants in the wavefunction grow as $\log(\log(2^{N}))$. Below we explicitly report the dominant excitations considered along with their respective coefficients (chosen to be $\leq 0.721$). The choice of determinants is based on excitations that remain close to the valence (occupied) orbitals. The determinants are reported using interleaved notation where occupancies are represented as $(\alpha,\beta,\alpha,\beta,\ldots)$:
  • For $N=4$:
    |$\ket{\Psi}=0.70\ket{1100}+0.71\ket{0011}.$
  • For $N=6$:
    |$\ket{\Psi}$ $=$ $0.60\ket{110000}-0.6\ket{001100}$
    $+$ $0.529\ket{000011}.$
  • For $N=8$:
    |$\ket{\Psi}$ $=$ $0.60\ket{11000000}-0.6\ket{00110000}$
    $+$ $0.529\ket{00001100}.$
  • For $N=10$:
    |$\ket{\Psi}$ $=$ $0.72\ket{1100000000}-0.6\ket{0011000000}$
    $+$ $0.245\ket{0000110000}-0.245\ket{1001000000}.$
  • For $N=12$:
    |$\ket{\Psi}$ $=$ $0.72\ket{110000000000}-0.6\ket{001100000000}$
    $+$ $0.245\ket{100100000000}-0.245\ket{011000000000}.$
  • For $N=14$:
    |$\ket{\Psi}$ $=$ $0.72\ket{11000000000000}-0.6\ket{00110000000000}$
    $+$ $0.245\ket{10010000000000}-0.245\ket{01100000000000}.$
  • For $N=16$:
    |$\ket{\Psi}$ $=$ $0.72\ket{1100000000000000}-0.6\ket{0011000000000000}$
    $+$ $0.245\ket{1001000000000000}-0.245\ket{0110000000000000}.$